OpenStudy (helpmeplease14):

A builder is remodeling a circular deck to increase the diameter from 8 feet to 14 feet. By how much will the circumference of the deck increase? By how much will the area of the deck increase?

OpenStudy (helpmeplease14):

@Mehek14

Mehek (mehek14):

ok so find the circumference of each and the area

OpenStudy (helpmeplease14):

Which formula for the circumference?

OpenStudy (helpmeplease14):

I think its C=pi2r

Mehek (mehek14):

circumference is 2πr

OpenStudy (kendricklamar2014):

\[Radius = Diameter / 2\]

OpenStudy (helpmeplease14):

K

OpenStudy (kendricklamar2014):

To find the increase, first we need to find the Circumference of the first deck

OpenStudy (kendricklamar2014):

\[C = 2*3.14*8 = 50.27..\]

OpenStudy (helpmeplease14):

A= 200.96

OpenStudy (kendricklamar2014):

Now we need to find the remodeled one:\[C = 2 * 3.14 * 14 = 87.96..\]

OpenStudy (kendricklamar2014):

Do you have any answer choices?

OpenStudy (kendricklamar2014):

@HelpMePlease14

OpenStudy (helpmeplease14):

No

OpenStudy (kendricklamar2014):

OK, so to find the increase we take:\[87.96 - 50.27 = 37.69 ; ~ Rounded: 38\]

OpenStudy (helpmeplease14):

I got 87.92

OpenStudy (helpmeplease14):

As the circumference with 14 inches

OpenStudy (kendricklamar2014):

For the Radius did you use 3.14 or pi

OpenStudy (helpmeplease14):

3.14

OpenStudy (kendricklamar2014):

Well its still 87 so... its not that important

OpenStudy (helpmeplease14):

C = 2 pi r C = 2 (3.14) 14 C = 6.28 (14) C = 87.92

OpenStudy (helpmeplease14):

Ok

OpenStudy (kendricklamar2014):

So the answer is 38(Rounded)

OpenStudy (helpmeplease14):

That's how much will the circumference of the deck increase?

OpenStudy (kendricklamar2014):

Yes that is how much the Circumference will increase

OpenStudy (helpmeplease14):

OK thx

OpenStudy (kendricklamar2014):

Your Welcome

OpenStudy (helpmeplease14):

How to simplify 414.48 into a whole number?

OpenStudy (kendricklamar2014):

Whole numbers: Numbers with no fractions, decimals, or negative. So we need to round 414.48

OpenStudy (helpmeplease14):

Will it be 414?

OpenStudy (helpmeplease14):

@KendrickLamar2014

OpenStudy (kendricklamar2014):

You got it =)

OpenStudy (helpmeplease14):

Thx

OpenStudy (helpmeplease14):

@KendrickLamar2014 can u help me with another problem?

OpenStudy (kendricklamar2014):

i'll try

OpenStudy (helpmeplease14):

K the question: How do circumference and area change when the radius of a circle is doubled? Give an example.

OpenStudy (helpmeplease14):

@KendrickLamar2014

OpenStudy (kendricklamar2014):

Didn't you already ask this Question: http://openstudy.com/study#/updates/552c3761e4b0d7047f5b0a11